OpenStudy (anonymous):

If i touch the end of an , lets say 1000V and 5 A and my overall body resistance is 1000ohms how can i determine the current and voltage flow through my body?

OpenStudy (anonymous):

just use the formula of ohm's law I =V/R

OpenStudy (shane_b):

The voltage part is simple enough: 1000V assuming you're grounded somewhere. For the current, you can use I=V/R as @gerryliyana said but that's assuming your body's resistance is constant with time which won't be the case. Having said that...I=V/R stills provide you a general idea of the current going through you.

OpenStudy (anonymous):

if i want to find the current, i will use I = 1000V/1000ohms but if i am going to find voltage pass through my body, I can still use the V=IR formula? like: V = 1000ohms x 5A or i have to use V = 1000ohms x 1A

OpenStudy (shane_b):

If the voltage is 1000V and the resistance of your body is 1000ohms, the current will be 1A. The amperage will be based on those two values since they are assumed to be constant. If however, you didn't know the voltage but did know the current and resistance, you could calculate the voltage using V = IR.